OpenStudy (anonymous):

If n-(1/n)=x, then what is n^2+(1/n^2) in terms of x?

OpenStudy (anonymous):

x^2-2 x^2-1 x^2 x^2+1 x^2+2

OpenStudy (anonymous):

Observe what happens when we square \(x\):$$n-\frac1n=x\\(n-1/n)^2=x^2\\x^2=n^2-2n\frac1n+\frac1{n^2}=n^2+\frac1{n^2}-2$$